# Transformer-based Wireless Symbol Detection Over Fading Channels

### v1

@@ VII Experiment
We present the experimental results and analyze the performances of DEFINED and IC-SSL in comparison with baseline algorithms. The results suggest that our model achieves superior performance not only with sufficient pilot data but also exhibits notable improvements in limited-data scenarios by effectively leveraging noisy feedback. Furthermore, DEFINED demonstrates robust performance in complex modulation tasks as well as mismatched channel distributions.
@@ VII-A Baseline Algorithms
We first describe several baseline algorithms, including the ICL method, the MMSE algorithm, and MMSE-DF which is a decision-feedback variant of MMSE.
@@ VII-A2 MMSE Algorithm
This is a coherent detection algorithm, which first estimates the channel with the assistance of pilots and then performs detections for the subsequent received signals using the estimated channel. Assuming a known pilot signal matrix $X=[x_{1},x_{2},\cdots,x_{k}]$, the received signal matrix is represented as $Y=HX+Z=[y_{1},y_{2},\cdots,y_{k}]$. In the case of Rayleigh fading, both the channel and noise follow complex Gaussian distributions, and the pair $(H,Y)$ is jointly Gaussian. The MMSE estimator for $H$ can be derived as: $\hat{H}_{k}^{\text{MMSE}}=YX^{\dagger}(XX^{\dagger}+\sigma^{2}I)^{-1}$. When the channel fading is not Gaussian, this is the linear MMSE (LMMSE) channel estimator. Then, with the $t$-th received signal $y_{t}$, the transmitted symbol $x_{t}$ is estimated by projection onto the closest symbol in $\mathcal{X}$:
@@ VII-A3 MMSE-DF Algorithm
MMSE-DF is an extension of the previous MMSE solution by sequentially using the decision feedback data as if they were new pilot pairs. Starting with $k$ pilots, we compute the MMSE estimator of $H$ and detect $\hat{x}_{k+1}$ using $y_{k+1}$, as described in Equation 6. The decision pair $(y_{k+1},\hat{x}_{k+1})$ is merged with the existing dataset and iteratively used to detect each subsequent signal until $\hat{x}_{T}$. The SER is plotted against the decision feedback-extended context sequence length.
@@ VII-B Experimental Results
During testing, we randomly sample 80,000 prompts to compute the mean SER across tasks. Channel realizations follow i.i.d. Rayleigh distributions in both training and testing (except in Section VII-B6), with the difference being that in the training phase, the dataset has different SNRs as described before. Results are presented for BPSK, QPSK, 16QAM, and 64QAM in the SISO system, and for BPSK and QPSK in a 2x2 MIMO system. We note that for methods that do not utilize feedback (such as ICL-ICL), the SER is plotted as a function of the pilot sequence length to evaluate the impact of different numbers of pilots. For methods that utilize decision feedback (such as ICL-DF, MMSE-DF, and DEFINED), SER is plotted as a function of the context sequence length. In this case, we emphasize that the adopted number of pilots is fixed, and we increase the context length by iteratively adding detected symbols. Specifically, the $t$-th point represents the SER of the estimator $\hat{x}_{t+1}$ using $t$ contexts, omitting the 0-th point. A glossary summarizing key training and testing configurations is provided in Table I for clarity.
@@
(a) SISO with BPSK, SNR = 5 dB, 1 Pilot, $\text{gain}_{\text{DF}}=11.2\%$.
(b) SISO with QPSK, SNR = 10 dB, 1 Pilot, $\text{gain}_{\text{DF}}=11.6\%$.
(c) SISO with 16QAM, SNR = 20 dB, 1 Pilot, $\text{gain}_{\text{DF}}=43.7\%$.
(d) SISO with 16QAM, SNR = 20 dB, 2 Pilots, $\text{gain}_{\text{DF}}=22.9\%$.
(e) SISO with 64QAM, SNR = 25 dB, 1 Pilot, $\text{gain}_{\text{DF}}=45.6\%$.
(f) SISO with 64QAM, SNR = 25 dB, 2 Pilot, $\text{gain}_{\text{DF}}=23.4\%$.
@@ VII-B1 Comparison with Baseline Algorithms under DF-testing
We first compare DEFINED with the baseline algorithms under DF-testing in scenarios with extremely limited pilot data: a single pilot in the SISO system and two pilots in the 2×2 MIMO system. The results are presented in Figure 9.
The DEFINED lines represent the performance of our proposed model during DF-testing, demonstrating a significant reduction in SER as more decision feedback data is incorporated, i.e., increased context sequence lengths. For instance, in Figure 9c, for 16QAM with one pilot, the SER decreases from 0.135 to 0.076 as the feedback context sequence length increases, achieving a $0.437$ SER gain. Remarkably, our model achieves performance comparable to conventional methods requiring over $4$ pilot pairs while using only a single pilot pair. This confirms that the Transformer effectively utilizes noisy feedback to improve detection performance with limited pilot data.
The ICL-DF line, which represents a Transformer trained with the ICL method but tested using decision feedback as if the detected symbols are pilots, performs significantly worse than our model, nearly coinciding with the MMSE-DF line. This observation highlights that models trained solely on clean data struggle with noisy feedback, as they simply treat all feedback data as clean. This underscores the importance of DF fine-tuning to effectively handle noisy feedback.
@@ VII-B2 Comparison with Baseline Algorithms Without Decision Feedback
We then compare our DEFINED model with baseline algorithms under ICL-testing (i.e., no DF). Complete results are grouped according to cases in SISO and MIMO systems under specific modulation schemes and SNR levels, as shown in Figure 10. The ICL-ICL line, representing the model trained and tested using the ICL method, demonstrates that with 30 pilots, the Transformer slightly outperforms the MMSE algorithm with 30 pilots during ICL testing. This improvement arises from the model’s ability to jointly perform channel estimation and symbol detection, leveraging the synergy between these tasks.
Importantly, DEFINED also performs well in ICL testing, as indicated by the DEFINED-ICL line, which closely aligns with the ICL-ICL line. This result suggests that ICL pre-training, followed by carefully designed loss functions during decision feedback fine-tuning, also enables the model to effectively learn from clean data. Additionally, together with the results from DF-testing, we demonstrate that a single pre-trained Transformer can perform both DF and no-DF tasks. Our DEFINED model adapts well to real-world symbol detection scenarios, excelling with ample pilot data and maintaining strong performance even with a single pilot.
(a) ICL-testing performance in the SISO system.
(b) ICL-testing performance in the MIMO system.
Figure 10: SER as a function of pilot sequence length under ICL-testing conditions, comparing DEFINED and vanilla ICL models in SISO and MIMO systems across various modulation schemes.
@@ VII-B3 Comparison with Different SNRs and Varying Pilot Lengths
At high SNR levels, reduced data noise enables more accurate detections from pilot data, enhancing DEFINED model performance and accentuating the downward SER trend. However, at very high SNRs, the already low initial SER limits further improvement. Our DEFINED model also performs robustly with minimal pilot data, including the extreme single-pilot cases. As pilot length increases, as shown in Figures 9d and 9f, all algorithms exhibit improved performance in DF inference.
@@ VII-B4 Comparison with Different Modulation Schemes
We conduct experiments using BPSK, QPSK, 16QAM, and 64QAM modulations in the SISO system, representing classification tasks with 2, 4, 16, and 64 classes, respectively. As modulation complexity increases, the detection task becomes more challenging, but we observe greater performance improvements with complex schemes, as reflected by a more pronounced SER decrease with additional feedback data in our DEFINED model.
Analysis of the Transformer’s output logit vector shows that nearly all of the incorrect detections occur within a small region around the ground-truth label in the constellation. This “typical error event” [38] suggests that even incorrect detections carry valuable information, as the noisy label is often near the correct one, enhancing the model’s detections. Thus, as modulation complexity increases, the compact constellation set allows noisy feedback to provide more useful information, leading to better SER gains with added feedback data.
These findings demonstrate that our DEFINED model effectively captures the constellation set’s geometry. It not only learns detections but also recognizes relationships between classes, often assigning higher probabilities to neighboring labels when errors occur. Due to inherent data noise – e.g., channel fading and additive noise – received signals with nearby latent labels in the constellation may overlap. As a result, the model sees adjacent labels as close neighbors, letting its detections retain valuable information, even when they are not entirely accurate.
@@ VII-B5 MIMO Results
The results for MIMO systems with BPSK and QPSK using two pilots are shown in Figure 11. Compared across different modulation schemes, SNR levels, our DEFINED model can still perform well for high-dimensional detection problems. In fact, DEFINED exhibits a more pronounced decrease in SER with the inclusion of additional decision feedback data, leveraging the underlying communication system structure more effectively.
(a) MIMO with BPSK, SNR = 10 dB, 2 Pilots, $\text{gain}_{\text{DF}}=16.8\%$.
(b) MIMO with QPSK, SNR = 15 dB, 2 Pilots, $\text{gain}_{\text{DF}}=38.6\%$.
Figure 11: SER as a function of context sequence length for MIMO systems with BPSK and QPSK under varying SNR levels and 2 pilots.
@@ VII-B6 Robustness to Channel Distribution Mismatch
We evaluate the robustness of our DEFINED model under channel distribution mismatches. Specifically, we consider a DEFINED model trained on a SISO 64QAM system with a single pilot and SNRs uniformly sampled from [20, 25] dB, consistent with the previous experiment.
First, the model is trained on a Rayleigh fading channel without a line-of-sight (LOS) component, $H_{\text{train}}\sim\mathcal{CN}(0,I)$, and tested on a Rician fading channel with an LOS component at SNR = 25 dB: $H_{\text{test}}\sim\sqrt{\frac{\kappa}{\kappa+1}}e^{j\theta}+\sqrt{\frac{1}{\kappa+1}}\mathcal{CN}(0,I),$ where $\kappa=4$ represents the LOS-to-NLOS power ratio, modeling suburban or rural environments [38], and $\theta$ is uniformly sampled to reflect varying LOS conditions. As shown in Figure 12a, DEFINED remains robust despite the channel mismatch, benefiting from the additional LOS component, which reduces the SER from 0.117 to 0.036.
Second, we evaluate a more challenging scenario where the model is tested in a Rayleigh fading channel with two pilots at SNR = 15 dB (which is below the range of SNRs in the training), introducing increased noise and pilot mismatch. As shown in Figure 12b, DEFINED achieves a $10.1\%$ SER gain, reducing the SER from 0.573 to 0.515. In this extremely noisy condition, MMSE-P30, which uses 30 pilots, achieves an SER of 0.484. This highlights the adaptability of DEFINED to varying pilots and noise levels, demonstrating its robustness under adverse conditions.
(a) SISO Rician 64QAM, SNR = 25 dB, 1 Pilot, $\text{gain}_{\text{DF}}=71.1\%$.
(b) SISO Rayleigh 64QAM, SNR = 15 dB, 2 Pilots, $\text{gain}_{\text{DF}}=10.1\%$.
@@ VII-C1 Maximum Likelihood Sequence Detection Algorithm
In our experiment, phase ambiguity is resolved by knowing the phase of the first symbol in the transmitted sequence at the receiver and searching for the optimal $S^{\text{MLSD}}$ within the subspace of $\mathcal{X}^{T}$. The average SER is then calculated for the remaining $(T-1)$ symbols, as the phase of the first symbol is known for a coherent interval of length $T$. Moreover, knowing the phase of the first symbol in SISO BPSK and QPSK scenarios is almost equivalent to having one pilot and is analogous to the setup used in previous baseline algorithms, which include one pilot and a total decision feedback enhanced context sequence of length $(T-1)$. This ensures a relatively fair comparison between the methods. Consequently, we enumerate $T$ from 2 to 31, compute the mean SER over tasks for detecting a sequence of $T$ received signals, and plot it at the index corresponding to a context sequence length of $(T-1)$ in the following results.
VII-C2 Results
(a) SISO with BPSK, SNR = 5 dB
(b) SISO with QPSK, SNR = 10 dB
Figure 13: Comparison of the DEFINED model with 1 and 2 pilots and the MLSD algorithm in SISO systems for BPSK and QPSK.
As shown in Figure 13, we compare the performance of the non-coherent MLSD algorithm with our DEFINED model in the SISO system using BPSK and QPSK modulations. The implementation of efficient MLSD algorithms has been an active area of research for decades [44]. However, since the problem is fundamentally NP-hard, computationally efficient algorithms often suffer from performance degradation (albeit small in some settings). We thus implement MLSD using an exhaustive search for varying sequence lengths. As the computational complexity grows exponentially with the sequence length, we only report the feasible results within some reasonable timeframe.
The DEFINED-P1 curves demonstrate that our model’s performance with one pilot under DF-testing matches that of the MLSD detector, which is the optimal sequence estimator with the additional first symbol phase information (almost equivalent to one pilot). Furthermore, with two pilots, our DEFINED model significantly outperforms the MLSD detector, as DEFINED directly minimizes the SER, as shown by the DEFINED-P2 line. The adaptability of DEFINED allows it to smoothly transition between non-coherent and coherent detection modes while consistently benefiting from decision feedback data.
@@ VII-C2 Results
VII-D IC-SSL Results
(a) IC-SSL SISO with 16QAM, SNR = 20 dB
(b) IC-SSL SISO with 64QAM, SNR = 25 dB
Figure 14: In-context semi-supervised learning in a SISO system with 16QAM and 64QAM.
We compare the performance of our IC-SSL model with the previous DEFINED model in Section V. As shown in Figure 14, for the relatively difficult modulation schemes (16QAM and 64QAM) in the SISO system, the IC-SSL model performs similarly or slightly better than the DEFINED model. This demonstrates that even without the pseudo-labels, the performance improvement over ICL remains consistent. It also shows that the Transformer jointly extracts information from both labeled data pairs and unlabeled features, gaining a better understanding of the feature space from the unlabeled data. Given the clustered and smooth nature of the feature distribution, the model can establish a more optimal classification decision boundary, leading to improved SER in subsequent decision-making, effectively performing SSL during inference.
On the other hand, the DEFINED model, which undergoes ICL pre-training followed by DF fine-tuning, ultimately achieves similar performance as the IC-SSL model in DF-testing. This reveals that the model acquires general capabilities akin to supervised learning during ICL pre-training, while DF fine-tuning effectively transitions the model from supervised to semi-supervised learning, even with the inclusion of pseudo-labels. This is enabled by the Transformer’s self-attention mechanism, which assigns varying weights to different parts of the input sequence for each output embedding, allowing the model to focus automatically on the most relevant information. We conjecture that during DF fine-tuning, the attention parameters are adjusted to prioritize the unlabeled features over the pseudo-labels, explaining the initial spike and gradual decrease in loss observed during fine-tuning, as shown in Figure 6.
Unlike previous studies on ICL, where prompts consist solely of feature-label pairs and Transformers are employed to learn mappings between them, our IC-SSL model generalizes the prompt by incorporating unlabeled features. This approach demonstrates that Transformers, beyond the conventional supervised learning paradigm, are remarkably powerful in extracting information and learning SSL mappings. This innovation significantly extends the applicability of Transformers beyond this particular wireless communication task.

### v2

@@ VII Experiment
[…]
We present the experimental results and analyze the performances of DEFINED in comparison with baseline algorithms. The results suggest that our model achieves superior performance not only with sufficient pilot data but also exhibits notable improvements in limited-data scenarios by effectively leveraging noisy feedback. Furthermore, DEFINED demonstrates robust performance in complex modulation tasks as well as mismatched channel distributions.
@@ VII-A Baseline Algorithms
We first describe several baseline algorithms against which DEFINED will be compared.
@@ VII-A2 MMSE Algorithm
This is the canonical coherent detection algorithm, which first estimates the channel using pilots and then performs detections for the subsequent received signals using the estimated channel. Assuming a known pilot signal matrix $X=[x_{1},x_{2},\cdots,x_{k}]$, the received signal matrix is represented as $Y=HX+Z=[y_{1},y_{2},\cdots,y_{k}]$. In the case of Rayleigh fading, both the channel and noise follow complex Gaussian distributions, and the pair $(H,Y)$ is jointly Gaussian. The MMSE estimator for $H$ can be derived as: $\hat{H}_{k}^{\text{MMSE}}=YX^{\dagger}(XX^{\dagger}+\sigma^{2}I)^{-1}$ for the i.i.d. fading case. Note that when the channel fading is not Gaussian, this is the linear MMSE (LMMSE) channel estimator. Then, with the $t$-th received signal $y_{t}$, the transmitted symbol $x_{t}$ is estimated by projection onto the closest symbol in $\mathcal{X}$:
@@ VII-A3 MMSE-DF Algorithm
MMSE-DF is an extension of the previous MMSE solution by sequentially using the decision feedback data as if they were new pilot pairs. This method belongs to decision-directed (or data-aided) channel estimation [42, 43]. Starting with $k$ pilots, we compute the MMSE estimator of $H$ and detect $\hat{x}_{k+1}$ using $y_{k+1}$, as described in Equation 6. The decision pair $(y_{k+1},\hat{x}_{k+1})$ is merged with the existing dataset and iteratively used to detect each subsequent signal until $\hat{x}_{T}$. The SER is plotted against the decision feedback-extended context sequence length.
@@ VII-B Experimental Results
During testing, we randomly sample 80,000 prompts to compute the average SER across tasks. Channel realizations follow i.i.d. Rayleigh distributions in both training and testing (except in Section VII-B6), with the difference being that in the training phase, the dataset has different SNRs as described in Section V-C1. Results are presented for BPSK, QPSK, 16QAM, and 64QAM in the SISO system, and for QPSK in a $2\times 2$ MIMO system. We note that for methods that do not utilize feedback (such as ICL-ICL), the SER is plotted as a function of the pilot sequence length to evaluate the impact of different numbers of pilots. For methods that utilize decision feedback (such as ICL-DF, MMSE-DF, and DEFINED), SER is plotted as a function of the context sequence length. In this case, we emphasize that the adopted number of pilots is fixed, and we increase the context length by iteratively adding detected symbols. Specifically, the $t$-th point represents the SER of the estimator $\hat{x}_{t+1}$ using $t$ contexts, omitting the 0-th point. A glossary summarizing key training and testing configurations is provided in Table I for clarity.
@@ VII-B1 Comparison with Baseline Algorithms under DF-testing
We first compare DEFINED with the baseline algorithms under DF-testing in a variety of scenarios, including some with the most extremely limited pilot data: a single pilot in the SISO system. The results are presented in Figure 8.
The DEFINED lines represent the performance of our proposed model during DF-testing, demonstrating a significant reduction in SER as more decision feedback data is incorporated, i.e., increased context sequence lengths. For instance, in Figure 8c, for 16QAM with a single pilot, the SER decreases from 0.0168 to 0.0075 as the feedback context sequence length increases, achieving a $55.3\%$ SER gain. Remarkably, our model achieves performance comparable to conventional methods that require more than $4$ pilots while DEFINED using only a single pilot. This confirms that DEFINED effectively utilizes noisy feedback to improve detection performance with limited pilot data. Furthermore, we note that increasing the amount of pilots from 1 to 2 would improve the detection accuracy for the early detected symbols due to more clean labels. As the context length increases, the performance gradually matches that of 1-pilot since the total (clean and pseudo) labels are roughly the same for both cases.
On the other hand, ICL-DF, which represents a Transformer trained with ICL method but deployed using decision feedback as if the detected symbols are pilots, performs significantly worse than DEFINED, nearly coinciding with MMSE-DF. This observation highlights that models trained solely on clean data struggle with noisy feedback, as they simply treat all feedback data as clean. This underscores the importance of DF fine-tuning to effectively handle noisy feedback.
@@ VII-B2 Comparison with Baseline Algorithms Without Decision Feedback
We then compare DEFINED with baseline algorithms under ICL-testing (i.e., no DF). Results are reported in Figure 9. The ICL-ICL result, representing the model trained and tested using the ICL method, demonstrates that with 30 pilots Transformer slightly outperforms the MMSE algorithm with 30 pilots during ICL testing. This improvement arises from the model’s ability to jointly perform channel estimation and symbol detection.
Importantly, DEFINED also performs well in ICL testing, as indicated by the DEFINED-ICL result, which closely aligns with the ICL-ICL line. This suggests that ICL pre-training, followed by carefully designed loss functions during decision feedback fine-tuning, empowers the model to effectively learn from clean data. Together with the results from DF-testing, we conclude that a single pre-trained Transformer can perform both DF and no-DF tasks. Our DEFINED model adapts well to real-world symbol detection scenarios, excelling with ample pilot data and maintaining strong performance even with a single pilot.
Figure 9: SER as a function of pilot sequence length under ICL-testing conditions, comparing DEFINED and vanilla ICL models in SISO systems across various modulation schemes.
@@ VII-B3 Comparison with Different SNRs and Varying Pilot Lengths
At high SNR levels, reduced data noise enables more accurate detections from pilot data, enhancing DEFINED model performance and accentuating the downward SER trend. However, at very high SNRs, the already low initial SER limits further improvement. Our DEFINED model still performs robustly with minimal pilot data, including the extreme single-pilot cases. As pilot length increases, as shown in Figure 8, all algorithms exhibit improved performance in DF inference.
@@ VII-B4 Comparison with Different Modulation Schemes
We conduct experiments using BPSK, QPSK, 16QAM, and 64QAM modulations, which correspond to classification tasks with 2, 4, 16, and 64 classes, respectively. As modulation complexity increases, the detection task becomes more challenging. However, somewhat counterintuitively, we observe larger performance gains with DEFINED: the SER decreases more sharply with additional feedback data (see Figs. 8).
Analysis of the Transformer’s output logit vector shows that nearly all of the incorrect detections occur within a small region around the ground-truth label in the constellation. This “typical error event” [32] suggests that even incorrect detections carry valuable information, as the noisy label is often very close to the correct one, which still provides reasonably accurate channel information. Thus, as modulation complexity increases, the compact constellation set allows noisy feedback to provide more useful information, leading to better SER gains with added feedback data.
These findings demonstrate that DEFINED effectively captures the constellation set’s geometry. It not only learns detections but also recognizes relationships between classes, often assigning higher probabilities to neighboring labels when errors occur. Due to inherent data noise, such as channel fading and additive noise, received signals with nearby latent labels in the constellation may overlap. As a result, the model sees adjacent labels as close neighbors, letting its detections retain valuable information, even when they are not entirely accurate.
Figure 10: SER as a function of context sequence length for MIMO systems with QPSK at SNR = 15 dB and 4 pilots ($\text{gain}_{\text{DF}}=50.2\%$).
@@ VII-B5 MIMO Results
The results for MIMO systems with QPSK using 4 pilots are given in Figure 10. Compared across different modulation schemes, SNR levels, DEFINED can still perform well for multi-dimensional detection problems. In fact, DEFINED exhibits a more pronounced decrease in SER with the inclusion of additional decision feedback data, leveraging the underlying communication system structure more effectively.
@@ VII-B6 Robustness to Channel Distribution Mismatch
As mentioned, handling training and testing channel distribution mismatch without ML model update is one of the most important benefits of ICL-based symbol detection, and we validate this feature in this subsection.
Specifically, we consider a DEFINED model trained on a SISO 64QAM system with a single pilot and SNRs uniformly sampled from [30, 40] dB, consistent with the previous experiment. First, the model is trained on a Rayleigh fading channel without a line-of-sight (LOS) component: $H_{\text{train}}\sim\mathcal{CN}(0,I)$, and tested on a Rician fading channel with an LOS component at SNR = 25 dB: $H_{\text{test}}\sim\sqrt{\frac{\kappa}{\kappa+1}}e^{j\theta}+\sqrt{\frac{1}{\kappa+1}}\mathcal{CN}(0,I),$ where $\kappa=4$ is the Ricean factor modeling suburban or rural environments [32], and $\theta$ is uniformly sampled to reflect varying LOS conditions. As shown in Figure 11a, DEFINED remains robust despite the channel mismatch, benefiting from the additional LOS component, which reduces the SER from $0.118$ to $0.038$. This result also validates the theoretical insight in Remark 1, that the Transformer exhibits inherent robustness to distributional shifts in $\mu_{0}$ and $\mu_{1}$, here manifested by the LOS perturbation in the Rician channel.
Second, we evaluate a more challenging scenario where the model is tested in a Rayleigh fading channel with two pilots at SNR = 25 dB (which is below the range of SNRs in the training), introducing increased noise and pilot mismatch. As shown in Figure 11b, DEFINED achieves a $23.2\%$ SER gain, reducing the SER from $0.138$ to $0.106$. In this extremely noisy condition, MMSE-P30, which uses 30 pilots, achieves an SER of $0.096$. This highlights the adaptability of DEFINED to varying pilots and noise levels, demonstrating its robustness under adverse conditions.
(a) SISO Rician 64QAM, SNR = 25 dB, 1 Pilot, $\text{gain}_{\text{DF}}=67.8\%$.
(b) SISO Rayleigh 64QAM, SNR = 25 dB, 2 Pilots, $\text{gain}_{\text{DF}}=23.2\%$.
@@ VII-C1 Maximum Likelihood Sequence Detection Algorithm
In our experiment, phase ambiguity is resolved by artificially injecting the phase of the first symbol in the transmitted sequence at the receiver and searching for the optimal $S^{\text{MLSD}}$ within the subspace of $\mathcal{X}^{T}$. The average SER is then calculated for the remaining $(T-1)$ symbols, as the phase of the first symbol is known for a coherent interval of length $T$. We note that this is a “genie” case that represents the best case scenario. Moreover, knowing the phase of the first symbol in SISO BPSK and QPSK scenarios is almost equivalent to having one pilot, and is analogous to the setup used in previous baseline algorithms that include one pilot and a total decision feedback enhanced context sequence of length $(T-1)$. This ensures a relatively fair comparison between the methods. Consequently, we enumerate $T$ from 2 to 31, compute the mean SER over tasks for detecting a sequence of $T$ received signals, and plot it at the index corresponding to a context sequence length of $(T-1)$ in the following results.
Figure 12: Comparison between DEFINED with one and two pilot pairs and the MLSD algorithm in SISO systems for BPSK at SNR = 15 dB and QPSK at SNR = 20 dB.
As shown in Figure 12, we compare the performance of the non-coherent MLSD algorithm with DEFINED in the SISO system using BPSK and QPSK modulations. The implementation of efficient MLSD algorithms has been an active area of research for decades [44]. However, since the problem is fundamentally NP-hard, computationally efficient algorithms often suffer from performance degradation (albeit small in some settings). We thus implement MLSD using an exhaustive search for varying sequence lengths. As the computational complexity grows exponentially with the sequence length, we only report the feasible results within some reasonable timeframe.
The DEFINED-P1 curves demonstrate that our model’s performance with one pilot under DF-testing matches that of the MLSD detector, which is the optimal sequence estimator with the additional first symbol phase information (almost equivalent to one pilot). Furthermore, with two pilots, DEFINED significantly outperforms the MLSD detector, as shown by the DEFINED-P2 line. The adaptability of DEFINED allows it to smoothly transition between non-coherent and coherent detection modes (via varying the amount of pilots) while consistently benefiting from decision feedback data.